Use properties of exponents to simplify :
a) (9^1/2+25^1/2)^2
b) (25^-1/3)^-3/2
c) (a^6)/27)^-2/3

Answer by stanbon(75887)

Use properties of exponents to simplify :
a) (9^1/2+25^1/2)^2
= (3*5)^2
= 15^2
= 225
-------------
b) (25^-1/3)^-3/2
= 25^(-1/2)
= 5^(-1)
= 1/5
------------
c) (a^6)/27)^-2/3
= (a^2/3)^(-2)
= (3/a^2)^2
= 9/a^4
==========
Cheers,
Stan H.